OpenStudy (anonymous):

(-1/8) (5/9) perform the indicated operations.

OpenStudy (anonymous):

-*+=- so the sign of the operation is negative 1*5=5 Numerator 8*9=72 Denominator thus, =-5/72

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

yes. this is in lowest terms?

OpenStudy (anonymous):

thanks.